# Concurrent Fermionic Simulation Gate

### v1

@@ III.2 Weak BPD
In the case of iswap the two columns (and rows) indicate the states $|01\rangle$ and $|10\rangle$ with $g$ being the interchanging strength between the two, which is nearly the interaction strength between the two qubits in the computational level, $\Delta$ the frequency detuning of the two qubits. Yet the very matrix of Eq.(11) can represent the cphase operator once the columns (and rows) are labelled by $|11\rangle$ and $|02\rangle$. Therefore $g$ for the cphase operator will be the interaction strength between the highest excited level in computational subspace of two qubits with the closest non-computational energy level.
Now consider that the state of two isolated qubits is evolved for a specific time $t$ only by the unitary operator $U(t)$, i.e. $U(t)=e^{-iHt}$. In the two-level picture we chose for the two gates in Eq.(11), the unitary state evolution operator can be represented as $U(t)=e^{-it\Delta/2}\left(\cos(\Omega t)\hat{1}-\ i\sin(\Omega t)\left(n_{z}\hat{Z}+n_{x}\hat{X}\right)\right)$, with $\hat{X}$ and $\hat{Z}$ being the Pauli matrices, the Rabi frequency $\Omega=\sqrt{(\Delta/2)^{2}+g^{2}}$ determining the rotation rate. The Pauli coefficients $n_{z}=-\Delta/2\Omega$ and $n_{x}=g/\Omega$ define the rotation axes.
@@
The cphase gate requires frequency detuning $\Delta$ between the two qubits. In this gate transition sends the photon outside of the computational level, however in order to minimize leakage, it is demanded that the final state goes back to $|11\rangle$ at the end of the cycle of two cphase evolutions. Therefore the total evolution adds a phase to the state $|11\rangle$.
@@
Note that this is not the only way to drive a tunable conditional phase. In another approach, one drives the $|10\rangle\leftrightarrow|02\rangle$ transition with zero detuning and stays at $|02\rangle$ for varying time to accumulate different conditional phases[39]. The accumulated phase is given by $\varphi=(\omega_{20}-\omega_{10}-\omega_{01})\tau$, with $\tau$ being the waiting time at $|20\rangle$. This gate can be implemented fast because $\omega_{20}-\omega_{10}-\omega_{01}$ is usually in the hundred MHz range. The total conditional phase combines the dynamical and geometric phases in the above cases. It is also possible to have a pure geometric conditional phase by properly choosing the evolution trajectory on the Bloch sphere[40].
III.3 Strong BPD
@@
Let us once again consider the case of two flux-tunable transmons coupled via a harmonic resonator, as depicted in the hardware schematic of Fig.(2), however ow we consider large Hilbert space for each transmon and coupler. Bichromatic parametric drives are applied to the device, with each drive acting on one transmon. For simplicity of analysis, we consider the setup where one drive is applied to each transmon individually. However, this is not a necessary condition from a theoretical perspective. Drives could instead be applied to the resonator, if it is also tunable, or both drives could be applied to a single transmon or the coupler.
@@
Neglecting the counter-rotating terms, we can write down the general Hamiltonian of the system:
@@
$\begin{split}&H=H_{0}+H_{d},\\ &H_{0}=\sum_{\alpha=1,2,c}\omega_{\alpha}\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}+\frac{\delta_{j}}{2}\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}(\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}-1)\\ &\quad\quad\quad+\sum_{j,k=\{1,2\}}g_{j}(\hat{a}^{\dagger}_{j}\hat{a}_{c}+H.c.),\\ &H_{d}=\sum_{j=1,2}\Omega_{j}\sin(\nu_{j}t+\phi^{d}_{j})\hat{a}^{\dagger}_{j}\hat{a}_{j}.\end{split}$ (15)
@@
where the subindex $c$ denotes the coupler, $g_{j}$ is the coupling strength between transmon $j$ and the coupler. For the case of a harmonic coupler one expects that $\delta_{c}=0$, but in general coupler can also be another qubit with non-zero anharmonicity $\delta_{c}$.
@@
The total Hamiltonian $H$ in Eq.(15) consists of the undriven Hamiltonian $H_{0}$ and the drive Hamiltonian $H_{d}$. The undriven Hamiltonian $H_{0}$ is the generalized Jaynes-Cummings. Hamiltonian[41]. The drive Hamiltonian has the standard form of parametric drives[42, 43, 44, 45, 46, 47, 48, 49, 37, 50].
@@
In order to accurately study the dynamics of the system, we first find a transformation $U_{0}$ that diagonalizes $H_{0}$, i.e. $\tilde{H}_{0}=U_{0}^{\dagger}H_{0}U_{0}$ with $\tilde{H}_{0}$ denoting diagonalized $H_{0}$. By transforming the drive Hamiltonian to the same diagonal frame where in $H_{0}$ is diagonalized, the drive Hamiltonian becomes $\tilde{H}_{d}=U_{0}^{\dagger}H_{d}U_{0}$, which is not necessarily a diagonal matrix. The total Hamiltonian in the diagonal frame is then $\tilde{H}_{0}+\tilde{H}_{d}$. Note that in this treatment, we are studying the dynamics in the dressed basis of the undriven Hamiltonian. Note that the diagonal transformation $U_{0}$ can be analytically found using the Bogoliubov transformation[51, 52, 53] or by exact numerical diagonalization. For a derivation using the Bogoliubov transformation, see Appendix (A).
Here, in order to perform precise analysis we numerically determine the diagonalization transformation matrix $U_{0}$ and compare results with what we find analytically in Appendix (A). It should be mentioned that this does not add too much computational cost when the undriven Hamiltonian $H_{0}$ is fixed.
@@
Let us denote dressed transmon states by $Q_{1}$ and $Q_{2}$, and dressed coupler states with $R$. In the Fock basis $\{|Q_{1}Q_{2}R\rangle\}$. Total Hamiltonian reads:
@@
For the full derivation and the expression of $\tilde{G}(t)_{Q_{1}Q_{2}R;Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}$ see Appendix (B).
III.3.1 Monochromatic parametric driving case
Let us consider the simple and special case of single parametric driving, namely ‘Monochromatic Paramateric Drive (MPD).’ We apply MPD to Q1 to evaluate the time-dependency of the coupling strength between the swap between transmons, $|010\rangle\leftrightarrow|100\rangle$. The drive frequency is resonant with the transition: $\nu_{1}=|\tilde{\omega}_{010}-\tilde{\omega}_{100}|$.
Figure 4: Static coupling strength, analytical and numerical. We show results from the exact diagonalization method and full numerical simulation. The analytical formula is valid for a wide range, from dispersive regimes to strong coupling regimes. 4(a) Static $\bar{g}_{100,010}$ vs qubit detuning. 4(b) Static $\bar{g}_{100,010}$ vs drive amplitude. 4(c) Static $\bar{g}_{110,020}$ vs qubit detuning. The dip around 300 MHz corresponds to the root of $J_{0}(x)+J_{2}(x)$. 4(d) Static $\bar{g}_{110,020}$ vs drive amplitude. The analytical formula can well capture non-linear dependence on $\Omega$.
In Appendix (B) we generalize the case where transmons and coupler are flux-tunable and we apply three parametric drives, two on qubits with frequencies $\nu_{1}$ and $\nu_{2}$, and one with frequency $\nu_{0}$ on the coupler. For this general case we worked out the resonant condition for transition $|Q_{1}Q_{2}R\rangle\leftrightarrow|Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}\rangle$ will be $\tilde{\omega}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}-\tilde{\omega}_{Q_{1}Q_{2}R}-(n_{0}\nu_{0}+n_{1}\nu_{1}+n_{2}\nu_{2})=0$ with integer numbers $n_{0}$, $n_{1}$, and $n_{2}$. The general resonance condition includes multi-photon processes, which indicates the presence of subharmonic interactions, see [54]. This condition for the case of MPD is simplified to the condition with $\nu_{0}=\nu_{2}=0$.
@@ III.3.1 Monochromatic parametric driving case
The coupling strength between $|110\rangle$ and $|020\rangle$ can be calculated similarly when the drive frequency $\nu_{1}$ is resonant with $\Delta_{110,020}=\tilde{\omega}_{110}-\tilde{\omega}_{020}$, see Appendix (B).
@@
    $\beta$: Reflects how differently the drives are coupled to the states $|100\rangle$ and $|010\rangle$, given by the difference of the diagonal terms $N^{1}_{100}-N^{1}_{010}$.
@@
    Drive amplitude $\Omega_{1}$ (numerator): Determines the drive strength. A stronger drive results in a more significant nonlinear effect.
@@
    Drive frequency $\nu_{1}$ (denominator): Indicates how rapidly the drive modulates the frequencies of the states. Faster modulation reduces the nonlinearity.
@@
The dips observed in Fig. (4 c) and Fig. (4 d) correspond to the condition $J_{2}\left(\beta\Omega_{1}/\nu_{1}\right)+J_{0}\left(\beta\Omega_{1}/\nu_{1}\right)=0$, which results in $g_{110,020}=0$. Solving the equation $J_{2}\left(\beta\Omega_{1}/\nu_{1}\right)+J_{0}\left(\beta\Omega_{1}/\nu_{1}\right)=0$ numerically yields the root $\beta\Omega_{1}/\nu_{1}\approx 3.83$, which matches well with the numerical simulation.
IV Crosstalk in BPD
Figure 5: Crosstalk effect of one drive on the other. (a) Numerical simulation of static $\bar{g}_{100,010}$ vs $\Omega_{2}$ and $\nu_{2}$ while the first drive is fixed at $\Omega_{1}=150$ MHz, $\nu_{1}=\Delta_{010,100}$, which is far detuned from $|010\rangle\leftrightarrow|100\rangle$ transition. Only the first drive is in resonance. The static coupling $\bar{g}_{100,010}$ is modulated by the second parametric drive. The dark stripes correspond to zeros of $J_{0}$ at $\Omega_{2}/\nu_{2}=2.40,5.52,8.65,\cdots$. The pattern is consistent with analytical calculation and Fig.(12). The anomalous resonance around $\nu_{2}=225$ MHz is not captured by RWA formula, because at this spot $2\nu_{2}=\Delta_{010,100}$ and 2-photon transition needs to be included to account for this resonance. (b) Numerical simulation of $P_{100}$ is consistent with analytical result except for 2-photon transition, see Fig.(12).
For the case of BPD with two parametric drives, the static coupling strength $g$ between two states can be evaluated similarly. For instance, the static coupling between $|010\rangle$ and $|100\rangle$ under one resonance drive $\Omega_{1}sin(\nu_{1}t)a_{1}^{\dagger}a_{1}$ and one off-resonant drive $\Omega_{2}sin(\nu_{2}t)a_{2}^{\dagger}a_{2}$ is given by:
$\bar{g}_{100,010}=i\alpha\Omega_{1}\left[J_{2}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)+J_{0}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)\right]J_{0}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)$ (21)
@@ IV Crosstalk in BPD
Details of derivation and even more general case of three parametric derives can be found in Appendix (B).
The first part of $\bar{g}{100,010}$, $i\alpha\Omega_{1}(J_{2}(\beta\Omega_{1}/\nu_{1})+J_{0}(\beta\Omega_{1}/\nu_{1}))$, is identical to the static $\bar{g}{100,010}$ derived under a single resonant drive, as shown in Eq. (20). The crosstalk effect of the off-resonant drive is introduced through the last term, $J_{0}(\gamma\Omega_{2}/\nu_{2})$. This term takes a similar form to the nonlinear term in the single-drive case. The crosstalk arises from the frequency modulation of the states $|100\rangle$ and $|010\rangle$ induced by the second drive.
@@
For small values of $\Omega_{2}/\nu_{2}$, $J_{0}$ is a monotonically decreasing function. As $\Omega_{2}/\nu_{2}$ increases, $J_{0}$ begins to oscillate around zero, resulting in multiple zeros in $\bar{g}{100,010}$ and $P{100}$. This behavior is illustrated in Fig. (5) and Fig. (12) in Appendix (C).
@@
Our result Eq.(23) also applies to multi-qubit systems. One can use it to suppress a target transition by requiring $\Pi_{k=2}^{N}J_{0}(\gamma_{k}\frac{\Omega_{k}}{\nu_{k}})=0$.
V Optimization and Results

### v2

@@ II.2 A Toy Model: Two Qutrits
In the case of iSWAP the two columns (and rows) indicate the states $|01\rangle$ and $|10\rangle$ with $g$ being the interchanging strength between the two, which is nearly the interaction strength between the two qubits in the computational level, $\Delta$ the frequency detuning of the two qubits. Yet the very matrix of Eq. (10) can represent the CPHASE operator once the columns (and rows) are labelled by $|11\rangle$ and $|02\rangle$. Therefore $g$ for the CPHASE operator will be the interaction strength between the highest excited level in the computational subspace of two qubits with the closest non-computational energy level.
Now consider that the state of two isolated qubits is evolved for a specific time $t$ only by the unitary operator $U(t)$, i.e., $U(t)=e^{-iHt}$. In the two-level picture we chose for the two gates in Eq. (10), the unitary state evolution operator can be represented as $U(t)=e^{-it\Delta/2}\left(\cos(\Omega t)\hat{1}-\ i\sin(\Omega t)\left(n_{z}\hat{Z}+n_{x}\hat{X}\right)\right)$, with $\hat{X}$ and $\hat{Z}$ being the Pauli matrices, the Rabi frequency $\Omega=\sqrt{(\Delta/2)^{2}+g^{2}}$ determining the rotation rate. The Pauli coefficients $n_{z}=-\Delta/2\Omega$ and $n_{x}=g/\Omega$ define the rotation axes.
@@
The CPHASE gate requires frequency detuning $\Delta$ between the two qubits. In this gate transition sends the photon outside of the computational level, however, in order to minimize leakage, it is demanded that the final state goes back to $|11\rangle$ at the end of the cycle of two CPHASE evolutions. Therefore the total evolution adds a phase to the state $|11\rangle$.
@@
This is not the only way to drive a tunable conditional phase. In another approach, one drives the $|10\rangle\leftrightarrow|02\rangle$ transition with zero detuning and stays at $|02\rangle$ for varying time to accumulate different conditional phases[51]. The accumulated phase is given by $\varphi=(\omega_{20}-\omega_{10}-\omega_{01})\tau$, with $\tau$ being the waiting time at $|20\rangle$. This gate can be implemented fast because $\omega_{20}-\omega_{10}-\omega_{01}$ is usually in the hundred MHz range. The total conditional phase combines the dynamical and geometric phases in the above cases. It is also possible to have a pure geometric conditional phase by properly choosing the evolution trajectory on the Bloch sphere[52].
II.3 Full Model: Transmon-Coupler-Transmon
@@
Let us once again consider the case of two flux-tunable transmons coupled via a harmonic resonator, as depicted in the hardware schematic of Fig. (1). Here we consider a larger Hilbert space for each transmon and coupler. Bichromatic parametric drives are applied to the device, with each drive acting on one transmon. For simplicity of analysis, we consider the setup where one drive is applied to each transmon individually. However, this is not a necessary condition from a theoretical perspective. Drives could instead be applied to the resonator, if it is also tunable, or both drives could be applied to a single transmon or the coupler[53].
As discussed at the beginning of this section, treating the full circuit Hamiltonian as in Eq. (2) and Eq. (3) is not an easy task, of which the complexity might obscure the physics behind BPD dynamics. We therefore decide to model the circuit as two Kerr nonlinear oscillators coupled via a harmonic oscillator. We furthermore assume that the drives only modulate the frequencies of the oscillators but not their anharmonicities. With those approximations, the dominant features of the physics are kept, and analytical formulas are still feasible.
@@
Neglecting the counter-rotating terms, we can write down the Hamiltonian of the system:
@@
$\begin{split}&H=H_{0}+H_{d},\\ &H_{0}=\sum_{\alpha=1,2,c}\omega_{\alpha}\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}+\frac{\delta_{j}}{2}\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}(\hat{a}^{\dagger}_{\alpha}\hat{a}_{\alpha}-1)\\ &\quad\quad\quad+\sum_{j,k=\{1,2\}}g_{jc}(\hat{a}^{\dagger}_{j}\hat{a}_{c}+H.c.),\\ &H_{d}=\sum_{j=1,2}\Omega_{j}\sin(\nu_{j}t+\phi^{d}_{j})\hat{a}^{\dagger}_{j}\hat{a}_{j}.\end{split}$ (14)
@@
where the subindex $c$ denotes the coupler, $g_{jc}$ is the coupling strength between transmon $j$ and the coupler. For the case of a harmonic coupler one expects that $\delta_{c}=0$, but in general the coupler can also have a non-zero anharmonicity.
@@
The total Hamiltonian $H$ in Eq. (14) consists of the undriven Hamiltonian $H_{0}$ and the drive Hamiltonian $H_{d}$. The undriven Hamiltonian $H_{0}$ is the generalized Jaynes-Cummings. Hamiltonian[54]. The drive Hamiltonian has the standard form of parametric drives[55, 56, 57, 58, 59, 60, 61, 62, 63, 48, 64].
@@
In order to accurately study the dynamics of the system, we first find a transformation $U_{0}$ that diagonalizes $H_{0}$, i.e. $\tilde{H}_{0}=U_{0}^{\dagger}H_{0}U_{0}$ with $\tilde{H}_{0}$ denoting diagonalized $H_{0}$. By transforming the drive Hamiltonian to the same diagonal frame where in $H_{0}$ is diagonalized, the drive Hamiltonian becomes $\tilde{H}_{d}=U_{0}^{\dagger}H_{d}U_{0}$, which is not necessarily a diagonal matrix. The total Hamiltonian in the diagonal frame is then $\tilde{H}_{0}+\tilde{H}_{d}$. Note that in this treatment, we are studying the dynamics in the dressed basis of the undriven Hamiltonian. Note that the diagonal transformation $U_{0}$ can be analytically found using the Bogoliubov transformation[65, 66, 67] or by exact numerical diagonalization. Here, in order to perform precise analysis we numerically determine the diagonalization transformation matrix $U_{0}$. This does not add too much computational cost when the undriven Hamiltonian $H_{0}$ is fixed.
@@
Let us denote dressed transmon states by $Q_{1}$ and $Q_{2}$, and dressed coupler(a harmonic resonator in our analysis ) states with $R$. In the Fock basis $\{|Q_{1}Q_{2}R\rangle\}$. Total Hamiltonian reads:
@@
For the full derivation and the expression of $\tilde{G}(t)_{Q_{1}Q_{2}R;Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}$ see Appendix (C).
II.3.1 Monochromatic parametric driving case
Let us consider the simple and special case of single parametric driving, namely ‘Monochromatic Parametric Drive (MPD).’ We apply MPD to Q1 to evaluate the time-dependency of the coupling strength between the swap between transmons, $|010\rangle\leftrightarrow|100\rangle$. The drive frequency is resonant with the transition: $\nu_{1}=|\tilde{\omega}_{010}-\tilde{\omega}_{100}|$.
Figure 3: Static coupling strength, analytical and numerical. Analytical results are obtained via Eq. (19). Numerical results are from solving time-depenent Schrödinger equation defined by Hamiltonians of Eq. (14), where the Hilbert space is truncated to 100 dimensions. 5 levels are used for each transmon, and 4 levels are used for the coupler mode. The analytical formula is valid for a wide range, from dispersive regimes to strong coupling regimes. 3(a) Static $\bar{g}_{100,010}$ vs qubit detuning. 3(b) Static $\bar{g}_{100,010}$ vs drive amplitude. 3(c) Static $\bar{g}_{110,020}$ vs qubit detuning. The dip around 300 MHz corresponds to the root of $J_{0}(x)+J_{2}(x)$. 3(d) Static $\bar{g}_{110,020}$ vs drive amplitude. The analytical formula can well capture non-linear dependence on $\Omega$.
In Appendix (C) we generalize the case where transmons and coupler are flux-tunable and we apply three parametric drives, two on qubits with frequencies $\nu_{1}$ and $\nu_{2}$, and one with frequency $\nu_{0}$ on the coupler. For this general case we worked out the resonant condition for transition $|Q_{1}Q_{2}R\rangle\leftrightarrow|Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}\rangle$ will be $\tilde{\omega}_{Q_{1}^{\prime}Q_{2}^{\prime}R^{\prime}}-\tilde{\omega}_{Q_{1}Q_{2}R}-(n_{0}\nu_{0}+n_{1}\nu_{1}+n_{2}\nu_{2})=0$ with integer numbers $n_{0}$, $n_{1}$, and $n_{2}$. The general resonance condition includes multi-photon processes, which indicates the presence of subharmonic interactions, see [68]. This condition for the case of MPD is simplified to the condition with $\nu_{0}=\nu_{2}=0$.
@@ II.3.1 Monochromatic parametric driving case
The coupling strength between $|110\rangle$ and $|020\rangle$ can be calculated similarly when the drive frequency $\nu_{1}$ is resonant with $\Delta_{110,020}=\tilde{\omega}_{110}-\tilde{\omega}_{020}$, see Appendix (C). This $g$ coupling is imaginary because of the rotating frame we chose. In this frame, the interaction resembles a Y-type rather than an X-type interaction.
@@
    $\beta$ reflects how differently the drives are coupled to the states $|100\rangle$ and $|010\rangle$, given by the difference of the diagonal terms $N^{1}_{100}-N^{1}_{010}$.
@@
    Amplitude $\Omega_{1}$ (numerator) determines the drive strength. A stronger drive results in a more significant nonlinear effect.
@@
    Frequency $\nu_{1}$ (denominator) indicates how rapidly the drive modulates the frequencies of the states. Faster modulation reduces the nonlinearity.
@@
The dips observed in Fig. (3c) and Fig. (3d) correspond to the condition $J_{2}\left(\beta\Omega_{1}/\nu_{1}\right)+J_{0}\left(\beta\Omega_{1}/\nu_{1}\right)=0$, which results in $g_{110,020}=0$. Solving the equation $J_{2}\left(\beta\Omega_{1}/\nu_{1}\right)+J_{0}\left(\beta\Omega_{1}/\nu_{1}\right)=0$ numerically yields the first root $\beta\Omega_{1}/\nu_{1}\approx 3.83$, which matches well with the numerical simulation. Other higher-order roots are beyond the accessible range.
II.3.2 Drive Crosstalk in BPD
Figure 4: Crosstalk effect of one drive on the other. Tthe first drive is fixed at $\Omega_{1}=150$ MHz, $\nu_{1}=\Delta_{010,100}$. The second drive is far detuned from $|010\rangle\leftrightarrow|100\rangle$ transition. We sweep drive amplitude $\Omega_{2}$ and frequency $\nu_{2}$ to see the crosstalk effects. 4(a) Numerical simulation of static $\bar{g}_{100,010}$. The static coupling $\bar{g}_{100,010}$ is modulated by the second parametric drive. The dark stripes correspond to zeros of $J_{0}$ at $\Omega_{2}/\nu_{2}=2.40,5.52,8.65,\cdots$. The anomalous resonance around $\nu_{2}=225$ MHz is not captured by RWA formula, because at this spot $2\nu_{2}=\Delta_{010,100}$ and 2-photon transitions need to be included to account for this resonance. 4(b) Numerical simulation of $P_{100}$ is consistent with analytical result except for 2-photon transition. 4(c) Analytical calculation of static $\bar{g}_{010,100}$ using Eq. (20). The analytical plot confirms numerical simulation in (a). 4(d) Analytical calculation of static $\bar{g}_{010,100}$. The x-axis is rescaled to illustrate the dependence of crosstalk effect on $\Omega_{2}/\nu_{2}$.
For the case of BPD with two parametric drives, the static coupling strength $g$ between two states can be evaluated similarly. For instance, the static coupling between $|010\rangle$ and $|100\rangle$ under one resonance drive $\Omega_{1}\sin(\nu_{1}t)a_{1}^{\dagger}a_{1}$ and one off-resonant drive $\Omega_{2}\sin(\nu_{2}t)a_{2}^{\dagger}a_{2}$ is given by:
$g_{1}=i\alpha\Omega_{1}\left[J_{2}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)+J_{0}\left(\frac{\beta\Omega_{1}}{\nu_{1}}\right)\right]J_{0}\left(\frac{\gamma\Omega_{2}}{\nu_{2}}\right)$ (20)
@@ II.3.2 Drive Crosstalk in BPD
Details of derivation and even more general case of three parametric derives can be found in Appendix (C).
The first part of $g_{1}$, $i\alpha\Omega_{1}(J_{2}(\beta\Omega_{1}/\nu_{1})+J_{0}(\beta\Omega_{1}/\nu_{1}))$, is identical to the static $\bar{g}_{100,010}$ derived under a single resonant drive, as shown in Eq. (19). The crosstalk effect of the off-resonant drive is introduced through the last term, $J_{0}(\gamma\Omega_{2}/\nu_{2})$. This term takes a similar form to the nonlinear term in the single-drive case. The crosstalk arises from the frequency modulation of the states $|100\rangle$ and $|010\rangle$ induced by the second drive.
@@
For small values of $\Omega_{2}/\nu_{2}$, $J_{0}$ is a monotonically decreasing function. As $\Omega_{2}/\nu_{2}$ increases, $J_{0}$ begins to oscillate around zero, resulting in multiple zeros in $\bar{g}_{100,010}$ and $P_{100}$. This behavior is illustrated in Fig. (4). We numerically compute the static coupling strength $\bar{g}_{1}$ and the population transfer and compare with the analytical results of Eq. (20). It can be clearly seen that the crosstalk is mainly coming through the factor $J_{0}(\gamma\Omega_{2}/\nu_{2})$.
@@
The formula Eq. (22) can be derived following the same line as Appendix (C). Here we sketch how it can be done. In the rotating frame, the coupling strength takes the form:$g\propto\Sigma_{j}\alpha_{j}\Omega_{j}\sin(\nu_{j}t)\exp\{i[\nu_{1}t+\Sigma_{k}\gamma_{k}\frac{\Omega_{k}}{\nu_{k}}\cos(\nu_{k}t)]\}$, where $\nu_{1}$ is the detuning of the targeted transition. One can then expand the exponents using the Jacobi–Anger identity and perform RWA. Because only drive 1 is resonant with the targeted transition and all other drives are off-resonant, the only terms left in the sum belong to the following two cases:
1. $j=1$ so $\sin(\nu_{1}t)$ cancels the phase factor $\exp(i\nu_{1}t)$
2. $j\neq 1$ but it cancels expansion terms from $\exp(i\gamma_{j}\frac{\Omega_{j}}{\nu_{j}}\cos(\nu_{j}t))$ and the phase factor $\exp(i\nu_{1}t)$ is canceled by one expansion term from $\exp(i\gamma_{1}\frac{\Omega_{1}}{\nu_{1}}\cos(\nu_{1}t))$.
Assuming $J_{0}(...)>>J_{1}(...),J_{2}(...)$, it is easy to find that only the first case gives dominant contribution. This thus gives the result Eq. (22), where we have renamed $\beta\equiv\gamma_{1}$. This expression also applies to multi-qubit systems. One can use it to suppress a target transition by requiring $\Pi_{k=2}^{N}J_{0}(\gamma_{k}\frac{\Omega_{k}}{\nu_{k}})=0$.
III Minimizing Leakage